INAF queue observing, 2013 Oct 31-Nov01

Observer: M. Fumana, R. Speziali
Telescope Operator: G. Becchetti
Support Astronomer: M. Edwards
Instrument: MODS

Summary

At the sunset clear sky.
LUCI not available for most of the night due to an electronic issue.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Fusi Pecci_07 07D Dual grating 30min  
Fusi Pecci_07 07E Dual grating 30min  
Fusi Pecci_07 07F Dual grating 30min  
Fusi Pecci_07 07G Dual grating 30min  
Fusi Pecci_07 07H Dual grating 30min  
Fusi Pecci_07 07I Dual grating 60min Bad seeing
Fusi Pecci_07 07L Dual grating 45min  
Vanzella_20 macsj0717 Dual grating 2h20min large dimm seeing, but gcs seeing ~1.1

Details

00:33 The TO opens the dome.

01:17 executeLUCIScript.sh campoA_mos_210_K_1.spec
  • Error: Instrument control: MOS panel
  • LUCI is not available. At the moment we don't know if it will be available tomorrow

01:24 Reconfigure for MODS

02:00 Slew to FusiPecci_07D * acqMODS B218_ls_PA85.acq * red 1st quadrant corrupted * through slit image is fine

02:11 execMODS B218_ls_05.obs
  • seeing ~1"

02:50 FusiPecci_07E: acqMODS B19_ls_264.acq

03:05 execMODS B19_ls_06.obs
  • seeing ~1"

03:42 acqMODS B193_ls.acq * Error in script eexcution * Comunication error: push updated button in MODS console and run the script again

04:00 FusiPecci_07F: execMODS B193_ls_07.obs

04:52 FusiPecci_07G: execMODS B224_ls_10.obs

05:30 FusiPecci_07H: acqMODS B179_ls.acq
  • faint guiding star (the only one available). We try to collimate on a brighter one

05:43 acqMODS B179_ls.acq

05:55 execMODS B179_ls_08.obs
  • seeing is raising ~1.2

06:33 FusiPecci_07I: acqMODS B311_ls_PA105.acq

06:45 execMODS B311_ls_09.obs
  • seeing ~ 1"
  • seeing ~1.8"

07:24 Because of the seeing, we run the previous script again execMODS B311_ls_09.obs
  • bad seeing ~2.5"
  • the last exposure has seeing <1.5"

08:04 FusiPecci_07L: acqMODS B386_ls_PA90.acq

08:06 Bad guide star, modify the script and send it again
  • seeing is now <1"

08:19 execMODS B386_ls_11.obs

09:15 Standard G191B2B

09:40 Vanzella_20A acquisition

09:40 mods.macsj0717_mos_1.obs mods1r 0064-0066 mods1b 0027-0029
  • seeing ~1"

11:07 mods.macsj0717_mos_2.obs mods1r 0067-0069 mods1b 0030-0032
  • seeing ~1"

12:15 mods.macsj0717_mos_2.obs mods1r 0070 mods1b 0031
  • aborted after 1 exposure. Twilight reaches at 12h12' - do not consider images #71(R) and #32(B)

00:50 Calibrations
  • ID523456 flats mods1r 0072-0073 mods1b 0035-0036
  • ID523456 lamps mods1r 0074==0076 mods1b 0037-0039

-- MarcoFumana - 01 Nov 2013
Topic revision: r3 - 05 Nov 2013, RobertoSpeziali
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback